# Geodetic Exploration of the Interior Structure of Enceladus

### v1

@@ 6.4 Feasibility of the required measurements
The static and time-varying gravity field of Enceladus can be recovered by tracking the motion of a spacecraft (Park et al., 2016, 2020, 2024). The shape and topography of Enceladus can be derived using spacecraft measurements such as imaging (Park et al., 2015, 2019, 2020) or altimetry (Smith et al., 2001). The accuracy of recovered parameters typically depends on the measurement sensitivity, orbiting geometry, and spacecraft operating conditions. With reasonable assumptions, we can compute the expected precision of recovered Enceladus gravity and tidal parameters calculated in Section 6.1 presented in Table 3 (Smith et al., 2001; Park et al., 2015, 2016, 2019, 2020, 2024).
Previously, Park et al. (2020) simulated a scenario where a spacecraft is in a stable orbit around Enceladus with approximately 60 inclination and periapsis and apoapsis altitudes of 150 km and 200 km, respectively. Assuming a typical Deep Space Network (DSN) X-band tracking capability, $k_{2}$ and $J_{3}$ could be easily recovered with accuracies of $<0.002$ and $0.4\times 10^{-6}$ (Park et al., 2020), respectively. Considering a similar orbit geometry, but with a reasonable mission duration and a more capable X-up/Ka-down tracking system, $k_{2}$ could be recovered with an accuracy of $<0.0002$ and $0.04\times 10^{-6}$ (Park et al., 2020), respectively.
If an imaging system with high resolution and accurate pointing capabilities is used, temporal changes in surface features can be measured with an accuracy better than approximately half a pixel using a stereo reconstruction technique (Park et al., 2020). Assuming an imaging system with 1-meter resolution and $<1$-meter pointing accuracy, the libration amplitude can be measured with $<1$-meter accuracy. Furthermore, tidal deformation (both radial and horizontal) can be measured with an accuracy $<50$ cm. This scenario assumes images of the same surface feature are collected under varying observation geometries and at different phases of Enceladus relative to Saturn.
An onboard altimeter would be ideal for directly measuring radial deformation (e.g., Steinbrügge et al., 2015; Thor et al., 2020). By assuming a similar orbiting geometry and sampling strategy with an altimetry precision of 1 m, the radial deformation could be recovered with an accuracy $<20$ cm using surface tie points (i.e. repeat observation of the same surface point at different orbital phase).
@@ 7 Conclusions and discussion
Determining the interior properties of Enceladus remains challenging given the currently available observations. We explored the possibility of determining the internal structure of Enceladus and the tidal heating processes by means of a suite of geodetic measurements. We presented constraints on the interior structure of Enceladus by revisiting the latest analysis of geodetic data obtained by Cassini. Based on these analysis, we provide estimates of the plausible ranges of the currently unavailable tidal response, i.e., $k_{2}$, $h_{2}$, and $l_{2}$, and their associated phase lags. The tidal Love numbers of Enceladus compatible with the currently available observations span a wide range depending on the interior properties. We showed that the potential measurement of the $k_{2}$ and phase lag can be explained by tidal heat generated in the shell or in the rocky core, presenting a case of degeneracy that requires further constraints to be resolved. Therefore, neglecting the part of the parameter space associated with core-dominated tidal dissipation can result in an erroneous interpretation of the time-variable gravity field for the interior structure. Our analysis shows that measuring tidal displacements and/or phase lags can discriminate between core-dominated and shell-dominated tidal dissipation cases. If $k_{2}$ is measured within 0.002, the measurement of the maximum vertical and horizontal displacements within 82 cm and 40 cm are sufficient to obtain a 2-$\sigma$ confidence for discriminating between the two heat generation scenarios. The phase lag associated with $h_{2}$ differs significantly between the two cases of tidal dissipation. In the case of a core-dominated tidal dissipation and an elastic ice shell, $\epsilon_{h_{2}}$ is negligible, while for a shell-dominated tidal dissipation, $\epsilon_{h_{2}}$ can reach tens of degrees (10–50). The requirements for measuring $h_{2}$ and $l_{2}$ are relaxed when improved constraints on $k_{2}$ are obtained. If $k_{2}$ is constrained within 0.0002, measuring the vertical and/or tangential tidal displacements within 116 cm and 51 cm, respectively, is sufficient to discriminate between the two classes of tidal dissipation. We note that while measuring the tidal phase lags is probably necessary to constrain the heat generation mechanisms in the case of Europa and Ganymede as considered by Hussmann et al. (2016), here we demonstrate that in the case of Enceladus, the anelastic deformations can be large enough to constrain the first-order heat generation scenario. Meanwhile, measurements of phase lags for Enceladus can considerably improve our constraints on the rheological properties of the interior.
We note that the computed values for shell and core viscosity should be considered as the “effective” viscosity. The viscosity profiles can depend on the variation of parameters such as porosity, temperature, and the water saturation percentile with depth. Nevertheless, even if other mechanisms than viscoelastic dissipation are producing the inferred tidal heating, the signature on the tidal response will remain the same.
@@
We also investigated the effect of the structural heterogeneities of the ice shell on the tidal response. Our results show that in the case of a largely elastic shell the surface displacements can significantly deviate from those of a spherically symmetric shell. These deviations can reach more than $\sim$50 cm for the thinnest parts of a heterogeneous shell compared to a spherically symmetric shell with the same global mean thickness. This observation is important for mission concept design purposes and for determining the locations at which the maximum displacements are measurable. We quantified the effect of a tidally active core on the tidal Love numbers of Enceladus with a heterogeneous ice shell. We found that if the core is tidally active to the extent needed to explain the steady-state heat budget, its effect on the observed $k_{2}$ in the presence of a heterogeneous shell can also be significant.
@@
We conducted inversions using synthetic observables and demonstrated their efficacy for probing the main interior structure parameters such as the rheology, density, and thickness of the layers, using a feasible set of measurements and uncertainties. As one of the important factors of determining habitability, we demonstrated that high- and low-density ocean scenarios differing by more than 40 psu (see Postberg et al., 2009b; Glein et al., 2018; Glein & Waite, 2020; Fifer et al., 2022; Postberg et al., 2023) can be distinguished using geodetic measurements. Such different interpretations originate mainly from the underlying assumptions on the provenance of the plume material, the potential change in their composition during their transport from the ocean, and the undetermined extent of overturning circulation and possible ice-ocean interactions (Zhu et al., 2017; Soderlund, 2019; Kang & Flierl, 2020; Lobo et al., 2021; Kang et al., 2022; Kang, 2022; Meyer et al., 2025; Ames et al., 2025; Bouffard et al., 2025). This obtained constraints on the density is only marginal, but it is strengthened when tighter constraints from the gravity field are obtained. Constraining core density and rheology can help distinguish between a low-density porous core and a higher-density rocky core and help us understand its composition, formation, and evolution.
@@
By extending the framework presented in this study to incorporate the future observations of total heat loss through thermal IR or microwave radiometry and comparison with the total tidally produced heat, we can understand whether Enceladus is in a thermal and orbital steady state, or alternatively, is experiencing a transient period during which the ocean may be becoming thicker or thinner. Such transient periods could occur in a cycle of freezing and melting due to a dynamically unstable orbit (Meyer & Wisdom, 2008) or in the case of an ongoing feedback between the moon’s thermal and orbital evolution of the body in the mean motion resonance (Ojakangas & Stevenson, 1986; Renaud et al., 2021). A transient ocean could also be the result of recent catastrophic events that affect the Saturnian system involving Enceladus (e.g., Asphaug & Reufer, 2013; Wisdom et al., 2022; Teodoro et al., 2023). By comparing the tidally generated heat in Enceladus with an independently measured surface heat loss rate, we can determine Enceladus’s thermal state and determine upper and lower bounds on the age of the subsurface ocean. The suite of needed geodetic measurements can be achieved by a future mission that provides additional constraints on the static gravity field, libration, tidal Love numbers, and rate of surface heat loss.
@@ 8 Acknowledgments
A. Bagheri acknowledges the support from the Swiss National Science Foundation by grant number P500PT_214435. A. Berne is supported by the Future Investigators in NASA Earth and Space Science and Technology (FINESST) Program (80NSSC22K1318).
@@ Appendix A Viscoelastic models
where $\Gamma(1+n)$ is the Gamma function and $\omega$ is the angular frequency. Here, $\alpha$ (frequency dependence) and $\beta$ are empirical parameters that are not well constrained. $\alpha$ is usually assumed to be in the range 0.1–0.4 based on laboratory experiments (e.g., Jackson & Faul, 2010) and $\beta$ is suggested to be approximately $\mu^{\alpha-1}/\eta^{\alpha}$ (Castillo-Rogez et al., 2011). Currently, these parameters are not well constrained for most planetary bodies (see e.g. Nimmo et al., 2012; Nimmo & Faul, 2013; Bagheri et al., 2019; Petricca et al., 2025). Given the current lack of such constraints for Enceladus, we follow Andrade (1962) and Faul & Jackson (2015), and use $\alpha=0.2$ and $\beta=1$. We investigate the effect of variation in the Andrade parameters in Figure 17. The approximation for $\beta\approx\mu^{\alpha-1}/\eta^{\alpha}$ (Castillo-Rogez et al., 2011) is employed. We find that the change in the real parts of the Love number as a function of the changes in the Andrade parameters within the range of conventionally assumed end-member values remains less than 10% if the dissipation is dominated in the shell and less than 20% if the dissipation is dominated in the core.
@@
$Q^{-1}_{B}=J_{2}(\omega)/J_{1}(\omega).$ (A6)
@@
Figure 17: Effect of changing the Andrade parameter $\alpha$ on the tidal response of Enceladus. Right (Left) panel, for a model of tidal dissipation dominantly in the shell (core).
Appendix B Derivation of tidal heating partitioning between the shell and the core
We can compute the tidal heating in the specific region of the satellite, such as the core or the shell, by integrating the radial shear dissipation in that part of the body as a result of tidal perturbations. Note that Equation (5) cannot be used for the core or the surrounding layers simply by replacing $k_{2}/Q$ or $\Im(k_{2})$ of the whole body by the tidal Love numbers of the core or the layers computed independently of the rest of the body. Here, we use equation 37 in Tobie et al. (2005) to derive the contribution of the core to the global tidal heating:
$H_{tidal}(r)=-\frac{21}{10}\frac{\omega^{5}e^{2}R_{E}^{4}}{r^{2}}H_{\mu}\Im{(\mu(r))},$ (B1)
@@ Appendix B Derivation of tidal heating partitioning between the shell and the core
[…]
Here, $\hat{k}_{2}^{core}$ is the piecewise gravitational tidal Love number associated with the core. For a homogeneous core, it can be computed from the deformation at the top of the core:
$\Im{(\hat{k}_{2}^{core}})=\frac{4\pi G\rho_{core}R_{core}^{4}}{5R_{E}^{3}}\Im{(\Delta R_{core})^{4}}$ (B4)
The tidal heating in the shell is computed using subtracting the contribution of the shell from the core:
$\dot{E}_{shell}=\dot{E}-\dot{E}_{core}.$ (B5)
@@
[…]
$\hskip 26.46085pt-\frac{21}{2G}\omega^{5}e^{2}R_{E}^{4}\bigg{(}R_{E}\Im{(k_{2})}-R_{core}\Im{(\hat{k_{2}}^{core})}\bigg{)}.$ (B6)
Equations (B2), (B5), and (B) provide a method to calculate the partitioning of global tidal heating for a given interior structure model. These equations require a priori information, assumptions, or independent constraints on the rheological and structural properties of the core to compute the shear dissipation in the core.
This method can be used in a forward model or in the context of an inverse problem based on a sampling approach. However, we consider requirements on the geophysical measurement that do not directly depend on the interior model parameters. Thus, we extend our analysis to derive the tidal heating on the basis of potentially observable quantities. Quantifying the partitioning of the heat in the icy body is equal to calculating the contribution of the core and the shell to the imaginary part of the gravitational Love number $\Im{(k_{2})}$, as indicated from Equation (5). Beuthe (2019) uses membrane mechanics to obtain the contribution of each layer within the body to the imaginary part of the tidal Love number:
$\Im{(k_{2})}=\left|\frac{k_{2}+1}{k_{2}^{o}+1}\right|^{2}\Im{(k_{2}^{o})}-\xi_{2}|h_{2}|^{2}\Im{(\Lambda_{T})},$ (B7)
@@
$\dot{E}_{\text{core}}=-\frac{5}{2}\frac{\omega R}{G}\left|\frac{k_{2}+1}{k_{2}^{0}+1}\right|^{2}\Im(k_{2}^{0})\langle|U_{2}^{(T)}|^{2}\rangle,$ (B8)
$\dot{E}_{\text{shell}}=\frac{5}{2}\frac{\omega R}{G}\xi_{2}\,|h_{2}|^{2}\Im(\Lambda_{T})\,\langle|U_{2}^{(T)}|^{2}\rangle,$ (B9)
@@
$\langle|U_{2}^{(T)}|^{2}\rangle=(\frac{21}{5})(\omega R)^{4}e^{2}.$ (B10)
The fluid crust gravitational Love number can be computed from Equation D.2 in Appendix D and Table 9 in Beuthe (2019). Once the tidal heating rate in the shell is computed, the partition of the heat produced in the core can be computed by subtracting the total heat from that of the shell. Following Beuthe (2015), the term $\Lambda_{T}$ in Equation (B8) is the complex spring constant of the shell and is computed as:
$\Lambda_{T}=\Lambda_{M}+\Lambda_{\rho}+\Lambda_{\chi}+\Lambda_{\omega}$ (B11)
@@
$\Lambda=f_{\mu}\hat{\mu}\epsilon,$ (B12)
$\Lambda_{\rho}=f_{\rho}\frac{\delta\rho}{\rho}\epsilon.$ (B13)
@@
$\nu=\frac{3K-2\mu}{6K+2\mu}.$ (B17)
This formulation provides a means for obtaining the tidal heating in the shell by measuring the tidal Love number $h_{2}$ (real and imaginary parts) and independent constraint on the imaginary part of the complex shear modulus of the shell, $\Lambda_{T}$, as shown in Equation (B8). Here, we extend the formulation to develop expressions for the partitioning of heat between the shell and the core, which do not require a priori knowledge of the shell rheology. Equation 125 in Beuthe (2015) provides an expression between the displacement and potential Love numbers, the rigidity of the shell, and the density structure of the body:
$k_{2}+1=(1+\Lambda_{T})h_{2}-5\frac{\delta\rho}{\rho}\epsilon.$ (B18)

### v3

@@ 6.4 Feasibility of the Required Measurements
The static and time-varying gravity field of Enceladus can be recovered by tracking the motion of a spacecraft (Park et al., 2016, 2020, 2024). The shape and topography of Enceladus can be derived using spacecraft measurements such as imaging (Park et al., 2019, 2020) or altimetry (Smith et al., 2001). The accuracy of recovered parameters typically depends on the measurement sensitivity, orbiting geometry, and spacecraft operating conditions. With reasonable assumptions, we can compute the expected precision of recovered Enceladus gravity and tidal parameters calculated in Section 6.1 presented in Table 3 (Smith et al., 2001; Park et al., 2016, 2019, 2020, 2024).
Previously, Park et al. (2020) simulated a scenario in which a spacecraft is in a stable orbit around Enceladus with approximately 60 inclination and periapsis and apoapsis altitudes of 150 km and 200 km, respectively. Considering a typical Deep Space Network (DSN) X-band tracking capability, $k_{2}$ and $J_{3}$ could be easily recovered with accuracies of <0.002<0.002 and $0.4\times 10^{-6}$ (Park et al., 2020), respectively. Considering a similar orbit geometry, but with a reasonable mission duration and a more capable Ka-up/Ka-down tracking system, $k_{2}$ could be recovered with an accuracy of <0.0002<0.0002 and $0.04\times 10^{-6}$ (Park et al., 2020), respectively.
If an imaging system with high resolution and accurate pointing capabilities is used, temporal changes in surface features can be measured with an accuracy better than approximately half a pixel using a stereo reconstruction technique (Park et al., 2020). Assuming an imaging system with 1-meter resolution and <1<1-meter pointing accuracy, the libration amplitude can be measured with <1<1-meter accuracy. This is partly possible because the reconstructed orbit accuracy using DSN data will be at the 1-m level. Furthermore, tidal deformation can be measured with an accuracy <50<50 cm (1-$\sigma$). This scenario assumes that images of the same surface feature are collected under varying observation geometries and at different phases of Enceladus relative to Saturn.
An onboard altimeter would be ideal for directly measuring radial deformation (e.g., Steinbrügge et al., 2015; Thor et al., 2020). By assuming a similar orbiting geometry and sampling strategy with an altimetry precision of <<5 m, the radial deformation could be recovered with an accuracy <25<25 cm (1-$\sigma$) using surface tie points (i.e. repeat observation of the same surface point at different orbital phases).
@@ 7 Conclusions and discussion
Determining the interior properties of Enceladus remains challenging given the currently available observations. We explored the possibility of determining the internal structure of Enceladus and the tidal heating processes by means of a suite of geodetic measurements. We presented constraints on the interior structure of Enceladus by revisiting the latest analysis of geodetic data obtained by Cassini. Based on these analyses, we provide estimates of the plausible ranges of the currently unavailable tidal response, i.e., $k_{2}$, $h_{2}$, and $l_{2}$, and their associated phase lags. The tidal Love numbers of Enceladus compatible with the currently available observations span a wide range depending on the interior properties. We showed that the potential measurement of the $k_{2}$ and phase lag can be explained by tidal heat generated in the shell or in the rocky core, representing a case of degeneracy that requires further constraints to be resolved. Therefore, neglecting the part of the parameter space associated with core-dominated tidal dissipation can result in an erroneous interpretation of the time-variable gravity field for the interior structure. Our analysis shows that measuring tidal displacements and/or phase lags can discriminate between core-dominated and shell-dominated tidal dissipation cases. If both real and imaginary parts of $k_{2}$ is measured within 0.002, the measurement of the maximum vertical and horizontal displacements within 82 cm and 40 cm are sufficient to obtain a 2-$\sigma$ confidence for discriminating between the two heat generation scenarios. The phase lag associated with $h_{2}$ differs significantly between the two cases of tidal dissipation. In the case of a core-dominated tidal dissipation and an elastic ice shell, $\epsilon_{h_{2}}$ is negligible, while for a shell-dominated tidal dissipation, $\epsilon_{h_{2}}$ can reach tens of degrees (10–50). The requirements for measuring $h_{2}$ and $l_{2}$ are relaxed when improved constraints on $k_{2}$ are obtained. If real and imaginary parts of $k_{2}$ is constrained within 0.0002, measuring the vertical and/or tangential tidal displacements within 116 cm and 51 cm, respectively, is sufficient to discriminate between the two classes of tidal dissipation. We note that while measuring the tidal phase lags is probably necessary to constrain the heat generation mechanisms in the case of Europa and Ganymede as considered by Hussmann et al. (2016), here we demonstrate that in the case of Enceladus, the anelastic deformations can be large enough to constrain the first-order heat generation scenario. Meanwhile, measurements of phase lags for Enceladus can considerably improve our constraints on the rheological properties of the interior.
@@
We also investigated the effect of the structural heterogeneities of the ice shell on the tidal response. Our results show that in the case of a largely elastic shell the surface displacements can significantly deviate from those of a shell of constant thickness. These deviations can reach more than $\sim$50 cm for the thinnest parts of a heterogeneous shell compared to a shell of constant thickness with the same global mean thickness. This observation is important for mission concept design purposes and for determining the locations at which the maximum displacements are measurable. We quantified the effect of a tidally active core on the tidal Love numbers of Enceladus with a heterogeneous ice shell. We found that if the core is tidally active to the extent needed to explain the steady-state heat budget, its effect on the observed $k_{2}$ in the presence of a heterogeneous shell can also be significant.
@@
We conducted inversions using synthetic observables and demonstrated their efficacy for probing the main interior structure parameters such as the rheology, density, and thickness of the layers, using a feasible set of measurements and uncertainties. As one of the important factors in determining habitability, we demonstrated that high- and low-density ocean scenarios differing by more than 40 psu (see Postberg et al., 2009; Glein et al., 2018; Glein & Waite, 2020; Fifer et al., 2022; Postberg et al., 2023) can be distinguished using geodetic measurements. Such different interpretations originate mainly from the underlying assumptions on the provenance of the plume material, the potential change in their composition during their transport from the ocean, and the undetermined extent of overturning circulation and possible ice-ocean interactions (Zhu et al., 2017; Soderlund, 2019; Kang & Flierl, 2020; Lobo et al., 2021; Kang, 2022; Meyer et al., 2025; Ames et al., 2025; Bouffard et al., 2025). This constraint obtained on the density is only marginal, but it is strengthened when tighter constraints from the gravity field are obtained. Constraining core density and rheology can help distinguish between a low-density porous core and a higher-density rocky core and help us understand its composition, formation, and evolution.
We highlight that the measurement requirements presented in this paper are computed under the assumption that the total produced heat produced is partitioned by at least a ratio of 90% and 10% between the core and the shell. These cases are the most likely scenarios, considering the patterns observed in Figure 1 where very limited values of ice and rock viscosities represent comparable dissipation in the shell and in the core. However, in a general case where partitioning of heat between the shell and the core are comparable, quantifying the tidal dissipation rate in the two parts of the body is possible by measuring the complex $k_{2}$ and $h_{2}$, and the density structure of the body (Equation 16). In this case, the uncertainty of quantifying the heat partitioning depends on the uncertainties on the parameters and observables involved in Equation (16). We also note that the computed values of shell and core viscosity should be considered as the “effective” viscosity. The viscosity profiles can depend on the variation of parameters such as porosity, temperature, and the water saturation percentile with depth. However, even if other mechanisms than viscoelastic dissipation are producing the inferred tidal heating, the signature on the tidal response will remain the same. We also point out that the methodology presented here does not address the possibility of tidal dissipation in the ocean, as a result of resonantly enhanced tidal flow as discussed in previous work (Hay & Matsuyama, 2017; Matsuyama et al., 2018; Tyler, 2020). A small $h_{2}$ amplitude and phase lag can rule out the occurrence of tidal dissipation in the shell, but may be associated with tidal dissipation occurring either in the ocean or in the core. However, significant ocean tidal heating at Enceladus is very unlikely to take place because it would require a very thin ocean (less than a few kilometers thick) to cause the resonance (Hay & Matsuyama, 2017; Matsuyama et al., 2018). Such a thin ocean is incompatible with the constraints we obtain on the interior structure from the available estimates of the libration, gravity field, and total mass of Enceladus.
@@
By extending the framework presented in this study to incorporate the future observations of total heat loss and comparison with the total tidally produced heat, we can understand whether Enceladus is in a thermal and orbital steady state, or alternatively, it is experiencing a transient period during which the ocean may become thicker or thinner. The rate of heat loss can be quantified through precise thermal IR or microwave radiometry, or combined constraints on the structure of the shell and its thermal conductivity. The orbital state can be studies by the ground-based or spacecraft astrometry. The suite of needed geodetic measurements can be achieved by a future mission that provides additional constraints on the static gravity field, topography, libration, tidal Love numbers, and rate of surface heat loss.
@@ 8 Acknowledgments
A. Bagheri acknowledges the support from the Swiss National Science Foundation by grant number P500PT_214435. A. Berne is supported by the Future Investigators in NASA Earth and Space Science and Technology (FINESST) Program (80NSSC22K1318). We also acknowledge the two anonymous reviewers for their constructive suggestions which helped improvement of the paper.
@@ Appendix A Viscoelastic models
where $\Gamma(1+n)$ is the Gamma function and $\omega$ is the angular frequency. Here, $\alpha$ (frequency dependence) and $\beta$ are empirical parameters that are not well constrained. $\alpha$ is usually assumed to be in the range 0.1–0.4 based on laboratory experiments (e.g., Jackson & Faul, 2010) and $\beta$ is suggested to be approximately $\mu^{\alpha-1}/\eta^{\alpha}$ (Castillo-Rogez et al., 2011). Currently, these parameters are not well constrained for most planetary bodies (see e.g. Nimmo & Faul, 2013; Bagheri et al., 2019; Bierson, 2024; Petricca et al., 2025). Given the current lack of such constraints for Enceladus, we follow Andrade (1962) and Faul & Jackson (2015), and use $\alpha=0.2$ and $\beta=1$. We investigate the effect of variation in the Andrade parameters in Figure 14. The approximation for $\beta\approx\mu^{\alpha-1}/\eta^{\alpha}$ (Castillo-Rogez et al., 2011) is employed. We find that the change in the real parts of the Love number as a function of the changes in the Andrade parameters within the range of conventionally assumed end-member values remains less than 10% if the dissipation is dominated in the shell and less than 20% if the dissipation is dominated in the core.
@@
$Q^{-1}_{B}=J_{2}(\omega)/J_{1}(\omega),$ (A6)
where $J_{1}$ and $J_{2}$ are the real and negative imaginary parts of the dynamic compliance $\hat{J}(\omega)$.
@@
Figure 14: Effect of changing the Andrade parameter $\alpha$ on the tidal response of Enceladus. Right and left panels show the tidal response parameters for models of tidal dissipation dominantly in the shell and in the core, respectively.
Appendix B Sensitivity of libration to elastic response
As highlighted in Section 4, if the ice shell is sufficiently thick as inferred for Enceladus, the libration amplitude is mostly determined by the moment of inertia and therefore, the thickness and density of the shell. We further explore the variations of the shell libration amplitude and the period of free libration as functions of the product of the shell density and shell thickness as a representative of its moment of inertia (Figure 15). We also show the same quantities with respect to the ratio of shell thickness to the shell density for better understanding of the behavior. As shown in the figure, the libration amplitude has dramatically large values for a certain value of $\rho_{shell}\times D_{shell}$ where the natural frequency of the shell equals the orbital period of Enceladus.
Figure 15: (a) Natural libration period (hr) and (b) the amplitude of forced physical libration of Enceladus as functions of the product of shell thickness and density as a representative of its shell MoI. (c,d) Same quantities as a function of the ratio of shell thickness and shell density. The plots are produced for fine grids of shell thickness and densities in ranges [1-60] km and [400-900] kg/m3. Colorbar indicates the associated shell thickness.
Elastic forces would be important only if the shell is very thin or its rigidity is very low. The effects of these parameters are also reflected in the natural (resonant) libration of the shell. We compare the effects of the shell rigidity against the shell thickness on the natural free libration period and the libration amplitude (Figure 16). As expected, libration is significantly more sensitive to the shell thickness compared to its rigidity. This different sensitivity is because for the shells that are not very thin compared to the radius of the body (such as those of Europa and Titan), the librational response is dominated by the inertial forces and the shell is effectively rigid. Note that as shown in Equation (8), two free libration frequencies exist for a body. Here we only show the natural frequency which is closer to the orbital period of Enceladus with the second free libration frequency is significantly longer for any reasonable shell properties and thus not relevant ($100~\rm hr$). Orbital period of Enceladus and the range of the inferred libration amplitudes has been demonstrated in the plots.
Figure 16: Natural libration period (hr) and the amplitude of forced physical libration of Enceladus as functions of the shell thickness and rigidity.
Because the rigidity of the shell does not play an important role in the librational response of the body, the viscosity of the shell has also a very small effect as well. Only in the case of a very low-viscosity shell with very low effective rigidity that can flexible respond to the tidal forces, the effect of the shell’s viscoelastic response becomes important (Figure 9b, panel a). As expected, the viscosity of the shell does not affect the libration even if highly tidally deformable.
Figure 17: Variation of the amplitude of the libration as a function of ice shell thickness and the effective viscosities of shell (a) and the core (b). In panel (a), the core viscosity is assumed equal to $\rm 10^{20}~Pa.s$. The region between the red lines shows the libration amplitude spanning the same range in Figure 3. Other interior structure parameters are the same as in Figure 2. At $\rm D_{shell}\approx 5$ km, the natural free libration period of the shell approaches the orbital period of Enceladus (for $\rho_{shell}$ fixed here to 900 kg/m3.). The shell thickness and amplitude of libration are plotted in logarithmic scales.
Appendix C Derivation of tidal heating partitioning between the shell and the core
We can compute the tidal heating in the specific region of the satellite, such as the core or the shell, by integrating the radial shear dissipation in that part of the body as a result of tidal perturbations. Note that Equation (5) cannot be used for the core or the surrounding layers simply by replacing $k_{2}/Q$ or $\Im(k_{2})$ of the whole body by the tidal Love numbers of the core or the layers computed independently of the rest of the body. Here, we use equations (33) and (37) in Tobie et al. (2005) to derive the contribution of the core to the global tidal heating:
$H_{tidal}(r)=-\frac{21}{10}\frac{\omega^{5}e^{2}R_{E}^{4}}{r^{2}}H_{\mu}\Im{(\mu(r))},$ (C1)
@@ Appendix C Derivation of tidal heating partitioning between the shell and the core
$\int_{0}^{R_{core}}H_{\mu}\Im{(\mu(r))dr}=\frac{5R_{core}}{4\pi G}\Im{(\hat{k}_{2}^{core}}).$ (C3)
Here, $\hat{k}_{2}^{core}$ is the piecewise gravitational tidal Love number associated with the core. For a homogeneous core, $k_{2}^{core}$ can be computed from the deformation at the top of the core (Van Hoolst et al., 2013; Shao & Nimmo, 2022, e.g.,):
$\Im{(\hat{k}_{2}^{core}})=\frac{4\pi G\rho_{core}R_{core}^{4}}{5R_{E}^{3}}\Im{(\Delta R_{core})^{4}}$ (C4)
The tidal heating in the shell is computed by subtracting the contribution of the core from the total heat:
$\dot{E}_{shell}=\dot{E}-\dot{E}_{core}.$ (C5)
@@
$\dot{E}_{shell}=-\frac{42\pi}{5}\omega^{5}e^{2}R_{E}^{4}\int_{R_{BO}}^{R_{E}}H_{\mu}\Im{(\mu(r))dr}$
$\hskip 26.17648pt=-\frac{42\pi}{5}\omega^{5}e^{2}R_{E}^{4}\bigg{(}\int_{0}^{R_{E}}H_{\mu}\Im{(\mu(r))dr}-\int_{0}^{R_{BO}}H_{\mu}\Im{(\mu(r))dr}\bigg{)}$
$\hskip 26.46085pt-\frac{21}{2G}\omega^{5}e^{2}R_{E}^{4}\bigg{(}R_{E}\Im{(k_{2})}-R_{core}\Im{(\hat{k_{2}}^{core})}\bigg{)}.$ (C6)
Here, $R_{E}$ and $R_{OB}$ refer to the radius at Enceladus’s surface at the bottom of the ocean. Equations (C2), (C5), and (C) provide a method to calculate the partitioning of global tidal heating for a given interior structure model. These equations require a priori information, assumptions, or independent constraints on the rheological and structural properties of the core to compute the shear dissipation in the core.
This method can be used in a forward model by assumptions on the core properties or in the context of an inverse problem based on a sampling approach. However, we consider requirements on the geophysical measurement that can not directly depend on the interior model parameters. Thus, we extend our analysis to derive the tidal heating on the basis of potentially observable quantities. Quantifying the partitioning of the heat in the icy body is equal to calculating the contribution of the core and the shell to the imaginary part of the gravitational Love number $\Im{(k_{2})}$, as indicated from Equation (5). Beuthe (2019) uses membrane mechanics to obtain the contribution of each layer within the body to the imaginary part of the tidal Love number:
$\Im{(k_{2})}=\left|\frac{k_{2}+1}{k_{2}^{o}+1}\right|^{2}\Im{(k_{2}^{o})}-\xi_{2}|h_{2}|^{2}\Im{(\Lambda_{T})},$ (C7)
@@
$\dot{E}_{\text{core}}=-\frac{5}{2}\frac{\omega R}{G}\left|\frac{k_{2}+1}{k_{2}^{0}+1}\right|^{2}\Im(k_{2}^{0})\langle|U_{2}^{(T)}|^{2}\rangle,$ (C8)
$\dot{E}_{\text{shell}}=\frac{5}{2}\frac{\omega R}{G}\xi_{2}\,|h_{2}|^{2}\Im(\Lambda_{T})\,\langle|U_{2}^{(T)}|^{2}\rangle,$ (C9)
@@
$\langle|U_{2}^{(T)}|^{2}\rangle=(\frac{21}{5})(\omega R)^{4}e^{2}.$ (C10)
The fluid crust gravitational Love number can be computed from Equation D.2 in Appendix D and Table 9 in Beuthe (2019). Once the tidal heating rate in the shell is computed, the heat produced in the core can be computed by subtracting the heating in the shell from the total heat. Following Beuthe (2015), the term $\Lambda_{T}$ in Equation (C8) is the complex spring constant of the shell and is computed as:
$\Lambda_{T}=\Lambda_{M}+\Lambda_{\rho}+\Lambda_{\chi}+\Lambda_{\omega}$ (C11)
@@
$\Lambda=f_{\mu}\hat{\mu}\epsilon,$ (C12)
$\Lambda_{\rho}=f_{\rho}\frac{\delta\rho}{\rho}\epsilon.$ (C13)
@@
$\nu=\frac{3K-2\mu}{6K+2\mu}.$ (C17)
This formulation provides a means for obtaining the tidal heating in the shell by measuring the tidal Love number $h_{2}$ (real and imaginary parts) and independent constraint on the imaginary part of the complex shear modulus of the shell, $\Lambda_{T}$, as shown in Equation (C8). Here, we extend the formulation to develop expressions for the partitioning of heat between the shell and the core, which do not require a priori knowledge of the shell rheology. Equation 125 in Beuthe (2015) provides an expression between the displacement and potential Love numbers, the rigidity of the shell, and the density structure of the body:
$k_{2}+1=(1+\Lambda_{T})h_{2}-5\frac{\delta\rho}{\rho}\epsilon.$ (C18)